OpenStudy (anonymous):

2^x=4^x+1 help and thanx

OpenStudy (whpalmer4):

\[2^x=4^x+1\] or \[2^x=4^{x+1}\]?

OpenStudy (anonymous):

second one

OpenStudy (whpalmer4):

Okay, if you remember \((ab)^n = a^nb^n\) and \(2*2=4\) we can rewrite the right hand side of the equation like this: \[2^x = (2*2)^{x+1} = 2^{x+1}*2^{x+1}\]Can you do it from there?

OpenStudy (anonymous):

i dont think so @whpalmer4

OpenStudy (whpalmer4):

Okay. \[2^x = 2^{x+1}*2^{x+1}\]Doesn't that imply that \[x = x+1+x+1\]? When we multiply exponentials with the same base, we add the exponents...all of these have the same base, so that means the sum of the exponents on the left equals the sum of the exponents on the right. Surely you can solve that last equation?